# Scalable registration of single quantum emitters within solid immersion lenses through femtosecond laser writing

### v1

@@ Abstract
Optically active silicon-vacancy (VSi) centers in silicon carbide (SiC) serve as qubits, interfacing spins via photons. This capability allows the encoding of photonic information within the spin state and facilitates on-demand readout, promising applications such as quantum memories. However, electron irradiation, a common technique for creating defects in SiC, lacks spatial selectivity, limiting scalability. We employed femtosecond laser writing within photonic structures to generate single (VSi) centers, registering them to photonic structures and enhancing optical collection efficiency by a factor of 4.5. Characterization of 28 laser-written defects centers in solid immersion lenses (SILs) showed distributions relative to the photonic structure’s center of 260 nm in the x-direction and 60 nm in the y-direction, with standard deviations of () nm and () nm, respectively. This method is scalable for developing integrated quantum devices using spin-photon interfaces in SiC.
I Introduction
@@
Color centers in the solid state, such as optically-active point defects and impurities, are among the most prominent systems for quantum technology [1]. Spin-photon interfaces associated to color centers in diamond [2, 3, 4], silicon carbide (SiC) [5, 6, 7, 8, 9] and silicon (Si) [10, 11], along with single rare-earth ion dopants in crystals [12], have been used in pioneering demonstrations of long-distance quantum networks. The nitrogen-vacancy center in diamond, as well as other defects in SiC [13, 14, 15, 16] and hexagonal boron nitride [17, 18, 19, 20], are used in a variety of quantum sensing applications, ranging from probing condensed matter problems and fundamental physics, to healthcare [21].
@@
Enhancing photon collection efficiency through tailored photonic structures is critical for optically-read-out spin qubits, as it directly impacts spin-photon interface efficiency in quantum networks and sensitivity in quantum sensing [22, 23]. This has been achieved for example with simple structures that minimize the effect of total internal reflection inside a high-index material, such as solid immersion lenses (SILs) [24, 25, 26, 27], or light guiding structures such as waveguides [28, 29, 30, 31]. The alternative is to employ a microcavity, such as a nanopillar [32, 33, 34, 35, 36] or photonic crystal [37, 38], to enhance light-matter interaction, maximize the fraction of emission into the coherent zero-phonon line, and increase optical extraction efficiency [39, 40].
@@
A significant challenge in the scalable integration of quantum emitters into photonic structures is their registration into the area within the structure that provides the most substantial optical enhancement. Depending on the specific structure, this region can be on the order of about in SILs, down to a few tens of nanometers for resonant structures with small mode volumes, such as photonic crystal cavities. Two main approaches have registered single quantum emitters to photonic structures. The first consists of mapping the position of pre-existing emitters against an array of markers and then fabricating the structure concerning the same marker array [41, 42, 43]. This procedure provides accuracy down to a few nanometers [41] and enables pre-selecting emitters with optimal properties for the specific application, but it is extremely time-consuming and not easily scalable into large arrays.
@@
A second possibility is to implant the ion species required to create the emitter into existing photonic structures, for example, through the use of focused ion beams [44, 45, 46, 47]. This technique is very effective, but can only be used to create shallow emitters (depths of m depth), as the lateral accuracy of implanted ion placement is similar in scale to the implantation depth due to increased straggle within the material. Furthermore, the higher ion energies required for deeper implantation typically induce more considerable lattice damage, which may degrade the quantum emitter properties [48, 49, 50].
@@
Recently, the generation of quantum emitters by laser-writing has received increasing attention. This technique exploits highly energetic carriers, created either by below-bandgap illumination with a high-power femtosecond laser [51, 52, 53, 54] or above-bandgap illumination with lower laser power [55], to initiate an avalanche ionization process. Above-bandgap (UV) illumination has been used to create quantum emitters in SiC nanophotonic structures [55], and femtosecond NIR laser pulses have been utilized to generate ensembles of quantum emitters in diamond nano-cavities [45].
@@
Here, we create single quantum emitters registered to SILs by direct below-bandgap femtosecond laser writing. Femtosecond laser writing enables us to create intrinsic point defects based on vacancies directly in the focal region of the SILs. Therefore, this method does not require any alignment to markers and can be fast and compatible with wafer-scale processing. The photonic structure itself can enhance the laser-writing field, potentially removing the need for any registration effort. By using the focusing effects of the lens, weaker laser pulse energies can be used to generate emitters. In contrast to ion implantation and UV illumination, direct below-bandgap femtosecond laser-writing enables the creation of single quantum emitters deeper into micron-scale structures, which are less sensitive to surface noise and typically feature better quantum coherence properties.
@@
We focus on quantum emitters in SiC, a material that uniquely combines spin-photon interfaces [5, 6, 8] with long spin coherence times [56, 7] with the possibility of integrating photonic [57, 58, 59, 28] and microelectronic functionalities [60, 61]. Our approach readily adapts to alternative materials, such as diamond [24, 42, 62].
II Experiment details
The experiment was conducted using commercial 4H-SiC material (Xiamen PowerWay©), with substrate and epilayer thickness and respectively, and residual n-doping level of . The material was diced into chips, and arrays of hemispherical SILs with nominal radius were fabricated using the grayscale hard-mask lithography process set out in [26]. It is worth noting that this method provides a key combination of scalability to large arrays with minimal time and cost increase and control over the shape of high-aspect-ratio microstructures, as opposed to other techniques which are either slow and expensive (e.g. focused ion beam milling [63] and two-photon polymerization [64, 65]) or provide limited shape control and aspect ratio (e.g. resist reflow lithography [66, 67, 68, 69], self-assembled microspheres [70], and microscale screen printing [71]).
The ability to create large arrays of shape-controlled lenses enables systematic parameter characterization when laser writing into SILs, an essential step due to the beam-focusing effects of the lens. This effect means that the parameters from writing through a planar interface cannot be straightforwardly applied within SILs, as described in the following section.
II.1 Laser writing procedure
@@
The laser writing process for generating quantum defects in the center of prefabricated SILs in 4H-SiC is illustrated in Fig. 1 a. The laser writing was performed using a regeneratively amplified Ti:Sapphire laser source, delivering 250 fs pulses at a wavelength of 790 nm and a repetition rate of 1 kHz (shown in Supplementary Figure S2). The laser was focused through a high-NA oil immersion objective lens (Olympus 60, 1.4 NA).
@@
A single femtosecond laser pulse with adjustable pulse energy was directed to selected focal points within the 4H-SiC sample. The degree of induced lattice damage can be controlled by varying the pulse energy, both in the case of writing in the planar region and within SILs. For planar interfaces, the focal depth was set to 5 m, matching the radius of the hemispherical monolithic SILs (depicted by the red spots in Fig. 1 a). Focusing into high-index materials like 4H-SiC induces strong spherical aberrations, which were corrected using a spatial light modulator (SLM).
@@
During fabrication, the laser was focused through the SIL to its central axis at a depth corresponding to the sample’s planar surface (represented by the blue spots in Fig. 1 a). This adjustment aimed to enable precise defect placement with minimal lattice perturbation and ensure optimal axial alignment between the SIL focus and laser induced defects.
@@
The laser writing process was initially characterized by photoluminescence (PL) under 532 nm CW optical excitation (1 mW) in a home-built confocal microscope. In both SIL and planar configurations, the intensity of the PL decreases with a lower pulse energy, as shown in Fig. 1 b. For each pulse energy level, data were collected from five spatially separated sites (in the case of SILs, 5 different neighboring SILs) fabricated in a single column. All data points represent the arithmetic mean of intensities measured across these five equivalent fabrication points. Saturating behavior is observed in both cases, where the PL intensities detected from defect centers stabilize with respect to the writing power in the regime of high pulse energy.
@@
Notably, significantly lower pulse energies are required for laser writing through the SIL interface compared to the planar interface. This is evidenced by the plotted curve shown in Fig. 1 b shifting to a lower pulse energy regime for the SIL interface (1.4 to 4.0 nJ) relative to the planar interface (4.0 to 20 nJ), highlighting the additional focusing effect provided by the SIL. Furthermore, the saturation count rate for the SIL interface reaches 600 kcps, compared to 300 kcps for the planar interface. The upward shift in the saturation curve indicates the enhanced light extraction efficiency achieved through the SIL interface.
II.2 Laser writing-induced defect generation mechanisms
Defect generation by ultrafast laser pulses in transparent materials is typically modeled by three mechanisms: multiphoton ionization (MPI), tunneling ionization (Zener breakdown), and avalanche ionization [72]. We take the model for generating the GR1 defect (neutral carbon vacancy) in diamond [73] as our starting point. The 250 fs laser pulses used here are shorter than the characteristic (nanosecond) timescales of thermal diffusion [74], allowing thermal effects to be neglected. The transition between MPI and tunneling ionization is described by the Keldysh parameter [72]. MPI dominates over tunneling ionization when
where kg is the effective mass of lattice electrons, and eV are the refractive index and bandgap of 4H-SiC and is the frequency of the photons. Substituting these values, the intensity threshold is calculated as . The pulse energy is then determined as , where the beam waist is nm (for the planar interface) and pulse duration is fs. Substituting these parameters for planar interface. For the SIL interface, the effective NA increases to around 2.6, which reduces the beam waist to nm, and the pulse energy is then calculated to be . The curvature of the SIL increased the effective NA , tightening the beam waist and intensifying the field strength in the focal volume, which ultimately lowers the pulse energy threshold.
For very high pulse energies, MPI does not dominate and additional energy may not create new defects due to saturation of the available defect sites. In contrast to the model in [73] for the GR1 center in diamond, we consider a saturation model alongside the power law:
where is the amplitude coefficient, is the power law exponent, and is the saturation parameter. Experimental results for the SIL interface yield and , while for the planar interface, and . The enhanced non-linearity in SIL compared to the planar interface may arise from a stronger Zener breakdown contribution at lower pulse energies, possibly alongside other processes such as above-threshold ionization. Fig. 1 b demonstrates excellent agreement () with data from the planar interface, showing PL intensity saturation around 8 nJ. This aligns with the Keldysh prediction of MPI dominance below 16 nJ. The photon energy for defect generation is times the photon energy from 790 nm laser, yielding eV for the planar interface and eV for the SIL interface, both of them are significantly lower than the displacement energies for Si (66 eV) and C (24 eV) in 4H-SiC, indicating defect formation mediated via laser-generated hot charge carriers rather than direct atomic displacement[75].
III Results
III.1 Characterisation of laser-written quantum emitters
@@
The laser-written defects were characterized at room temperature with a home-built confocal microscope, as described in our earlier work [26]. A 780 nm diode laser provided continuous-wave optical excitation, and the emitted light passed through several long-pass filters at 800 nm before PL detection. By spatially mapping this PL emission over SILs that were laser-written into, we could identify laser-induced defects localized near the center of SILs. (Fig. 2).
@@
For each SIL where a laser-written spot was detected, we performed measurements to determine the excitation power and brightness at which the emission saturates (power saturation measurement) and the normalized depth of the second-order correlation function at zero delay (), using low-jitter (average 40ps for the two channels) superconducting nanowire single-photon detectors (SNSPD, Single Quantum EOS). There is noticeable background luminescence from the surface (and edge of the SIL) but not within the SIL itself. This agrees with previous similar studies in 4H-SiC that notice intrinsic surface defects related to the surface oxide [55].
@@
Stochastic distributions of emitting defects consistent with Poissonian statistics could be detected across the written region, with the number and brightness of defects increasing with increasing laser-writing power. To minimize the creation of multiple emitters, writing must ideally be performed in the regime where the expectation value of creating any emitters is much less than one; for example, for a Poissonian expectation value of , the probability of creating a single emitter is , while the probability of creating two or more emitters is . So, even for such a low yield, two emitters would be created approximately 5% of the time. In the case of the region with laser writing pulse energy PE=1.8 nJ shown in Figure 2, 9 written defects occur within a set of 30 SILs, including five confirmed single photon emitters and two defects consistent with two emitters, leading to a Poissonian expectation value of approximately .
@@
The resulting values are reported in Table 1; we show the second-order correlation results before and after background subtraction. Fabrication using laser pulse energy 1.8 nJ created the most amount of emitters with , which is also illustrated in Figure 2. The SIL ID in Table 1 is with respect to the column and row of the SIL on the sample. For example, T30 is on column T and row 30 of the SILs processed, the characterisation for the spot in this SIL is also shown in Figure 3. The photoluminescence counts at saturation are also shown for each of these emitters.
We classify each PL spot based on the value to be either a single emitter (), multiple emitters , or as a PL spot not in the single-emitter regime . Therefore, as we decrease the laser writing power, we expect the likelihood of generating an emitter with () to increase so that there is an optimal laser writing power to generate single emitters reliably. This shows how narrow the window for effective fabrication of single color centers is using laser writing in SILs structures.
@@
To confirm the nature of the created quantum emitters, we examined their PL spectrum at low-temperature to identify characteristic zero-phonon lines (ZPL). The sample was placed in a cryostat at T= 4 K (Montana s100 Cryostation) and probed with a custom confocal setup (see Cilibrizzi et al. [8]). The emitters are excited with a 780 nm CW laser, and the PL spectra are measured with a grating spectrometer (OceanOptics QE Pro)). All retrieved spectra are shown in Supplementary Figure S5. Remarkably, we observe a wide spread of ZPL wavelengths in the spectral region 858nm-985nm. Out of 39 emission defects examined, 6 can have spectral lines that can be classified as silicon-vacancy (VSi) centers; these were, in particular, 1 V1’ center (ZPL at 858 nm), 4 V1 centers (h-site, ZPL at 861 nm) and 1 V2 center (k-site, ZPL at 916 nm). The remaining defects exhibit ZPLs at different wavelengths across this 858-985 nm range. These lines do not appear to match other unidentified lines previously reported in this wavelength range [76, 77, 78], but do fall within the theoretically-predicted range of VSi centers modified by nearby carbon anti-sites [79] and experimentally-observed range for silicon vacancies in etched membranes [80].
@@
We further performed optically-detected magnetic resonance (ODMR) measurements at room temperature, in the frequency ranges associated with V1 (zero-field splitting 4 MHz) and V2 (zero-field splitting 70 MHz) centers. For as-written emission defects, no ODMR signals were observed in these ranges. However, after annealing the sample at 600∘C for 30 minutes in vacuum (mbar), a process known to improve the yield of VSi centers, we observed an ODMR signal associated with V2 for a single SIL (of 12 measured), though with a large linewidth of about MHz (see Fig S6 in Supplementary Information).
@@
One possible explanation for the broad linewidths in the PL spectra and the lack of ODMR signal might be that the femtosecond laser creates other defects that are not optically active, in addition to the VSi. It is for example known, from molecular dynamics simulations [75] and photoluminescence and deep-level transient spectroscopy (DLTS) experiments [81, 82, 83], that carbon atoms feature much lower displacement energy than silicon atoms in SiC, so that the creation of e.g. carbon vacancies (V) is favored. The presence of charge traps, related to carbon vacancies and other defects [84, 60, 61], is expected to create electric and magnetic noise that may broaden the optical and magnetic resonance linewidths, respectively. Further experiments, beyond the scope of this manuscript, are needed to clarify the physics of femtosecond laser generation of point defects in SiC and optimize the laser writing parameters to improve the quality of generated quantum emitters.
III.2 Properties of single laser-written within a SIL
@@
In Figure 3, the key properties are given of a single color center generated in SIL T30 with a laser writing pulse energy . The photoluminescence map (Fig. 3 a) shows a spot centered well with respect to the SIL profile (dark circle). The photoluminescent background outside of the SIL region corresponds to surface defects on the etched surface of the SiC, postulated to arise at the SiC-surface oxide interface [55, 28, 85, 86, 87]. Spectroscopy of the emitted light at cryogenic temperatures (T = 4K) yielded a ZPL at 861 nm, consistent with a V1-type VSi center. Furthermore, even without background subtraction, the normalized second-order correlation at zero delay , confirms that the emission comes from one single photon emitter. Taking power saturation measurements of the single VSi center in SIL T30 and a single electron-irradiated VSi center beneath a planar interface (Fig. 3, blue and green, respectively), the performance of the SIL-registered emitter could be benchmarked. Following our previous work [26], we determine an optical collection efficiency enhancement factor of 4.5 ( vs. ) and a power intensification factor of 9.1 ( vs. ), consistent with the upper levels of performance observed for single VSi centers generated randomly throughout the SIL by electron irradiation.
III.3 Registration of emitters to SILs
@@
As mentioned earlier in Section II.1, laser writing through the SIL interface was carried out by manually aligning the laser focus to the center of the SIL at a depth of m, equivalent to the height of the lens structure. Contributions to the final emitter misalignment with respect to this target could be expected both from the random probability of generating a defect within the volume of highest writing intensity and the error in manual alignment, however these could be reduced by the focusing effect of the lens.
@@
To quantify these distributions, spatial PL maps of SILs with emitter spots were characterized, using image analysis software (ImageJ) to extract the center of each SIL and emitter spot (further discussed in the Supplementary Information ). The absolute displacements of all measured emitters from their mean position are given in Fig. 4 a. The statistics follow a Rayleigh distribution with scale parameter m (dashed line, for reference).
@@
When the absolute positions of the emitters with respect to the SIL are plotted (Fig. 4 b), clustering near the SIL center is evident, especially compared to emitter generation through electron irradiation in similar structures [26]. Focusing on the center region of the SIL (Fig. 4 c), it is apparent that the standard deviation of positions in the x-direction (nm) is higher than in the y-direction (nm). This is possibly due to the raster protocol for laser writing, where the writing proceeds along all columns in a row first before moving to the next row, so more steps are taken in x than in y.
Aligning to the SIL center can prevent the average position of generated defects from being at the center of the SIL due to systematic offsets from positioning or the writing angle. This is in contrast to writing in planar interfaces, where the locations of generated emitters are compared to the array of writing locations so that no mean offset is expected. This is also observed in Figure 4 c, where the mean emitter position is offset by 260 nm in x and 60 nm in y from the SIL center.
IV Conclusions

### v2

@@ Abstract
The precise registration of solid-state quantum emitters to photonic structures is a major technological challenge for fundamental research (e.g. in cavity quantum electrodynamics) and applications to quantum technology. Standard approaches include the complex multi-step fabrication of photonic structures on pre-existing emitters, both registered within a grid of lithographically-defined markers.
Here, we demonstrate a marker-free, femtosecond laser writing technique to generate individual quantum emitters within photonic structures. Characterization of 28 defect centers, laser-written at the centers of pre-existing solid immersion lens structures, showed offsets relative to the photonic structure’s center of 260 nm in the x-direction and 60 nm in the y-direction, with standard deviations of nm and nm, respectively, resulting in an average 4.5 times enhancement of the optical collection efficiency. This method is scalable for developing integrated quantum devices using spin-photon interfaces in silicon carbide and easily extendable to other materials.
Keywords: Color Centres, Femtosecond Laser Writing, Silicon Carbide, Scalable Quantum Devices, Qubits
@@
Color centers in the solid state, such as optically-active point defects and impurities, are among the most prominent systems for quantum technology [1]. Spin-photon interfaces associated to color centers in diamond [2, 3, 4], silicon carbide (SiC) [5, 6, 7, 8, 9] and silicon [10, 11], along with single rare-earth ion dopants in crystals [12], have been used in pioneering demonstrations of long-distance quantum networks. The nitrogen-vacancy center in diamond, as well as other defects in SiC [13, 14, 15, 16] and hexagonal boron nitride [17, 18, 19, 20], are used in a variety of quantum sensing applications, ranging from fundamental physics to healthcare [21].
@@
Enhancing photon collection efficiency through tailored photonic structures is critical for optically-read-out spin qubits, as it directly impacts spin-photon interface efficiency in quantum networks and sensitivity in quantum sensing [22, 23]. This has been achieved, for example, with simple structures that minimize the effect of total internal reflection inside a high-index material, such as solid immersion lenses (SILs) [24, 25, 26, 27, 28], or light guiding structures such as waveguides [29, 30, 31, 32, 33, 34]. The alternative is to employ a microcavity, such as a nanopillar [35, 36, 37, 38, 39], bullseye cavity [40] or photonic crystal [41, 42], to enhance light-matter interaction, maximize the fraction of emission into the coherent zero-phonon line, and increase optical extraction efficiency [43, 44].
@@
A significant challenge in scaling the integration of quantum emitters into photonic structures is registering them to the location that provides the greatest optical enhancement. Depending on the structure, this region can range from in SILs to for photonic crystal cavities. One approach to registering single quantum emitters in photonic structures consists of mapping the position of pre-existing emitters against a marker array, which can be aligned to during photonic structure fabrication [45, 46, 47]. This procedure could provide accuracy down to a few nanometers [45] and enables pre-selection of emitters with optimal properties for integration. However, it is extremely time-consuming and not easily scalable to large arrays.
@@
A second possibility is to implant the ion species required to create the emitter into existing photonic structures, for example, through the use of focused ion beams [48, 40, 49, 50]. This technique is very effective, but can only be used to create shallow emitters (depths of m depth), as the lateral accuracy of implanted ion placement is similar to the implantation depth. Higher-energy ions for deep implantation can degrade emitter quality [51, 52, 53].
@@
Recently, the generation of quantum emitters by laser-writing has received increasing attention. This technique exploits highly energetic carriers, created either by below- [54, 55, 56, 57, 33, 58] or above-bandgap [59] illumination with a high-power femtosecond laser to initiate an avalanche ionization process. Above-bandgap illumination has been used to create quantum emitters in SiC nanophotonic structures [59], and femtosecond NIR laser pulses have been utilized to generate ensembles of quantum emitters in diamond nano-cavities [40].
@@
Here, we create single quantum emitters registered to SILs by direct below-bandgap femtosecond laser writing. Femtosecond laser writing enables us to create intrinsic point defects based on vacancies directly in the focal region of the SILs. This method does not require alignment to markers and can be fast, making it compatible with wafer-scale processing. The photonic structure itself can enhance the laser-writing field, potentially removing the need for any registration effort. By using the focusing effects of the lens, weaker laser pulse energies can be used to generate emitters. In contrast to ion implantation and above-bandgap illumination, direct below-bandgap femtosecond laser writing enables the creation of single quantum emitters deeper within micron-scale structures, where they are less sensitive to surface noise and typically exhibit better quantum coherence properties.
@@
We focus on quantum emitters in SiC, a material that uniquely combines spin-photon interfaces [5, 6, 8] possessing long spin coherence times [60, 7] with integrated photonic [61, 62, 63, 29] and microelectronic [64, 65] functionalities. Our approach readily adapts to alternative materials, such as diamond [24, 46, 66].
The experiment was conducted using commercial 4H-SiC material (Xiamen PowerWay©) diced into chips (see SI section 1 for details). Arrays of hemispherical SILs with nominal radius were fabricated on these chips using the grayscale hard-mask lithography process set out in our previous work [26].
@@
The laser writing process for generating quantum defects in the center of prefabricated SILs in 4H-SiC is illustrated in Fig. 1 a. The laser writing was performed using a regeneratively-amplified Ti:Sapphire laser source, delivering 250 fs pulses at a wavelength of 790 nm and a repetition rate of 1 kHz (shown in Fig. S1 and S2 in SI Section 2). The laser was focused through a high-NA oil immersion objective lens (Olympus 60, 1.4 NA).
@@
A single femtosecond laser pulse with adjustable pulse energy was focused within the 4H-SiC sample. The degree of induced lattice damage can be controlled by varying the pulse energy, both when writing in the planar region and within SILs. For planar interfaces (red spots in Fig. 1 a), the focal depth was set to 5 m, matching the radius of the hemispherical monolithic SILs. Focusing into high-index materials like 4H-SiC induces strong spherical aberrations, which were corrected using a spatial light modulator (SLM).
@@
During writing, the laser was focused through the SIL along its central axis at a depth corresponding to the sample’s planar surface (blue spots in Fig. 1 a). This adjustment aimed to enable precise defect placement with minimal lattice perturbation and ensure optimal axial alignment between the SIL focus and laser-induced defects.
@@
The laser writing process was initially characterized by photoluminescence (PL) under 532 nm CW optical excitation (1 mW) in a home-built confocal microscope with a nm collection window (details discussed in SI section 2). In both SIL and planar configurations, the PL intensity increases with higher pulse energy, as shown in Fig. 1 b. For each pulse energy, data were collected from five spatially-separated sites (in the SIL case, 5 different neighboring SILs) written in a single column. All data points represent the arithmetic mean of intensities measured across these five equivalent fabrication points. Saturating behavior is observed in both cases, where the PL intensities detected from defect centers stabilize with respect to the writing power in the high pulse energy regime.
@@
Notably, significantly lower pulse energies are required for laser writing through the SIL compared to the planar interface, evidenced by the lower pulse energy range of the blue (SIL) curve in Fig. 1b (1.4 to 4.0 nJ) relative to the red (planar interface) curve (4.0 to 20 nJ). This highlights the additional focusing effect provided by the SIL. The larger error bars in PL counts through the SIL interface arises from minor deviations in SIL geometry and alignment uncertainties during focusing the writing laser, both of which introduce aberrations to the incident wavefront and associated variation in focal intensity. Furthermore, the saturation count rate for the SIL interface reaches 600 kcps, compared to 300 kcps for the planar interface, which indicates the light extraction efficiency achieved through the SIL interface is also enhanced.
To elucidate the underlying mechanisms of laser-induced defect generation, we consider established models for ultrafast laser interactions in transparent materials, which typically involve multiphoton ionization (MPI), tunneling ionization (Zener breakdown), and avalanche ionization [67]. As a starting point, we adopt the framework developed for the creation of GR1 centers (neutral carbon vacancies) in diamond [68], which shares key similarities with the processes in 4H-SiC. The 250 fs laser pulses employed in this work are significantly shorter than the characteristic timescales of thermal diffusion (nanoseconds) [69], allowing us to neglect thermal effects and focus on non-thermal ionization dynamics. The relative contributions of MPI and tunneling ionization are governed by the Keldysh parameter [67], with MPI dominating when the laser intensity satisfies the condition:
where kg is the effective mass of lattice electrons, and eV are the refractive index and bandgap of 4H-SiC, and is the frequency of the photons. Substituting these values, the intensity threshold is calculated as . The pulse energy is then determined as , where for the planar interface the beam waist is nm and pulse duration is fs. Substituting these parameters, . For the SIL interface, the effective NA increases to , which reduces the beam waist to nm, and the pulse energy is then calculated to be . At very high pulse energies, MPI loses dominance, resulting in a deviation from the expected power-law scaling, as shown in Fig. 1 b. This potentially leads to lattice breakdown which is associated with the onset of broadband PL emission and a saturation plateau in the photoluminescence count rate, as observed in Fig. 1b. Assuming that the PL intensity relates to the number of defects generated by the laser pulse; we fit this saturation model to the PL intensity vs. pulse energy data in Fig. 1b (solid curves) to a saturation model:
where is the amplitude coefficient, is the power law exponent, and is the saturation parameter (in slight contrast to the model presented in [68] for the negatively charged carbon vacancy center in diamond). Fitting for the SIL interface yields and , while for the planar interface, and . The enhanced non-linearity in the SIL compared to the planar interface may arise from a stronger Zener breakdown contribution at lower pulse energies. The defect generation process involves photon energies that are multiples of the 790 nm laser photon energy. For the planar interface, this results in an energy of eV, while for the SIL interface, it yields eV. These energies exceed the SiC bandgap, enabling the generation of hot charge carriers rather than direct lattice disruption, given the non-polar covalent bonding in SiC. Instead, multiphoton absorption excites carriers which transfer energy to the lattice, leading to defect formation. This carrier-mediated mechanism resembles Frenkel defect formation in diamond [70], highlighting a shared pathway for laser-induced structural modification in wide-bandgap materials.
@@
PL measurements were performed at room temperature using a 780nm laser, with 800nm dichroic and shortpass filters (Thorlabs FES800) placed in the excitation arm to suppress residual laser light, as described in previous work [26]. By spatially mapping this PL emission over laser-written SILs, we could identify laser-induced defects localized near the center of SILs (Fig. 2).
@@
For each SIL where a laser-written spot was detected, we measured the excitation power and brightness at which the emission saturates (power saturation measurement) and the normalized depth of the second-order correlation function at zero delay (), using low-jitter ( ps) superconducting nanowire single-photon detectors (Single Quantum EOS). There is noticeable background luminescence from the SiC/air interface but not within the SIL itself. This agrees with previous studies in 4H-SiC that report intrinsic defects related to native surface oxide [59].
@@
Stochastic distributions of emitting defects consistent with Poissonian statistics could be detected across the written region, with the number and brightness of defects increasing with increasing laser-writing power. To minimize the creation of multiple emitters, writing is ideally performed in the regime where the Poissonian expectation value of creating any emitters is much less than one (). For example, for an expectation value of , the probability of creating a single emitter is , while the probability of creating two or more emitters is . So, even for such a low yield, two emitters would be created approximately 5% of the time. In the case of the region with laser writing pulse energy PE=1.8 nJ shown in Fig. 2, 9 written defects occur within a set of 30 SILs, including five confirmed single photon emitters and two defects consistent with two emitters, leading to a Poissonian expectation value of .
@@
The resulting values are reported in Table 1; we show the second-order correlation results before and after background subtraction. Fabrication using laser pulse energy 1.8 nJ created the largest number of emitters with , see Fig. 2. The SIL ID in Table 1 denotes its grid location on the sample. We also show photoluminescence counts at saturation for each emitter.
We classify each PL spot based on its value to be either a single emitter (), multiple emitters , or as a PL spot not in the single-emitter regime .
@@
We examined the PL spectra of generated quantum emitters at low temperature (T= 4 K, Montana s100 Cryostation) to identify characteristic zero-phonon lines (ZPLs). The emitters were excited with a 780 nm CW laser, and the PL spectra were measured with a grating spectrometer (OceanOptics QE Pro)). For further information on the setup and measured spectra, see SI section 5.
@@
Remarkably, we observe a wide spread of ZPL wavelengths in the spectral region 858 nm-985 nm. Out of 39 emission defects examined, 6 have spectral lines that can be classified as silicon-vacancy (V) centers; these were 1 V1’ center (ZPL 858 nm), 4 V1 centers (h-site, ZPL 861 nm), and 1 V2 center (k-site, ZPL 916 nm). The remaining defects exhibit ZPLs at different wavelengths which do not appear to match other unidentified lines previously reported in this spectral range [71, 72, 73], but fall within the theoretically-predicted range of V centers modified by nearby carbon anti-sites [74] and experimentally-observed range for V centers in etched membranes [75].
@@
We further performed optically-detected magnetic resonance (ODMR) measurements at room temperature, in the frequency ranges associated with V1 (zero-field splitting 4 MHz) and V2 (zero-field splitting 70 MHz) centers. For as-written emission defects, no ODMR signals were observed in these ranges. However, after annealing the sample at 600∘C for 30 minutes in vacuum ( mbar), a process known to improve V yield, we observed an ODMR signal associated with V2 for one SIL (of 12 measured), though with a large linewidth of about MHz (see SI section 6 Fig. S6).
@@
One possible explanation for the lack of ODMR signal might be that the femtosecond laser creates other defects that are not optically active, in addition to the V. For example, carbon atoms feature much lower displacement energy than silicon atoms in SiC [76, 77, 78, 79], so that the creation of carbon vacancies (V) is favored. The presence of charge traps related to carbon vacancies and other defects [80, 64, 65] is expected to create electromagnetic noise that may broaden observed optical and magnetic resonance linewidths. Further experiments, beyond the scope of this work, are needed to clarify the physics of femtosecond laser generation of point defects and to optimize laser writing parameters for improving the quality of generated quantum emitters.
@@
In Fig. 3, the key properties are given of a single V center generated in SIL T30 with a laser writing pulse energy . The photoluminescence map (Fig. 3 a) shows a spot centered well with respect to the SIL profile (dark circle). The photoluminescent background outside of the SIL region corresponds to surface defects on the etched surface of the SiC, postulated to arise at the SiC-surface oxide interface [59, 29, 81, 82, 83]. Spectroscopy of the emitted light at cryogenic temperatures (T = 4K) yielded a ZPL at 861 nm, consistent with a V1-type V center. Furthermore, even without background subtraction, the normalized second-order correlation at zero delay , confirms that the emission comes from one single photon emitter. Taking power saturation measurements of the single V center in SIL T30 and a single electron-irradiated V center beneath a planar interface (Fig. 3, blue and green, respectively), the performance of the SIL-registered emitter could be benchmarked. Following our previous work [26], we determine an optical collection efficiency enhancement factor of 4.5 ( vs. ) and a power intensification factor of 9.1 ( and ), consistent with the upper levels of performance observed for single V centers generated randomly throughout the SIL by electron irradiation.
@@
Laser writing through the SIL interface was performed by microscope-aided alignment of the laser focus to the center of the SIL at a depth of m, equivalent to the height of the lens structure. Contributions to the final emitter misalignment with respect to this target could arise both from the random probability of generating a defect within the volume of highest writing intensity and the error in writing laser alignment.
@@
To quantify these distributions, spatial PL maps of SILs with emitter spots were characterized using image analysis software (ImageJ) to extract the center of each SIL and emitter spot (see SI section 3). The absolute displacements of measured emitters from their collective mean position are given in Fig. 4 a. The statistics follow a Rayleigh distribution with scale parameter m (dashed line).
@@
When the emitters’ absolute positions with respect to the SIL are plotted (Fig. 4 b), clustering near the SIL center is evident, especially compared to emitter generation through electron irradiation in similar structures [26]. Focusing on the SIL center (Fig. 4 c), we find that the standard deviation of positions in the x-direction ( nm) is higher than in the y-direction ( nm). This is possibly due to the raster protocol for laser writing, where writing proceeds along all columns in a row before moving to the next row, so that more steps are taken in x than in y.
When aligning the writing laser to SILs, systematic offsets in positioning or beam angle can lead to the mean defect position deviating from the structure’s center. This is observed in Fig. 4 c, where the mean emitter position is offset by 260 nm in x and 60 nm in y from the SIL center.